OpenStudy (anonymous):

Please help? I'm so confused. :\ "Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. a) passes through (-1, 4) and (-5, 2) b) slope 2, y-intercept -4 c) has an x-intercept of 6 and a y-intercept of 3 d) passes through (1, 2) with slope -5/3

OpenStudy (skullpatrol):

You need to start with definitions of what are slope-intercept, point-slope, and standard forms, right @Breezehh ???

OpenStudy (anonymous):

Yeah..

OpenStudy (anonymous):

a ) slope point form y - y1 = m(x - x1) where m = (y2 - y1)/(x2 - x1) b ) slope intercept y = mx + c c) intercept form a is x intercept and b is y x/a + y/b = 1 d)slope point form y - y1 = m(x - x1)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thank you. I don't understand it but I have to turn something in lol

OpenStudy (skullpatrol):

Askaway :)

OpenStudy (skullpatrol):

If you want to learn something, you have to put in the effort of learning.

OpenStudy (skullpatrol):

$$\Huge \text{a) passes through (-1, 4) and (-5, 2)}$$ $$\Huge \text {let }{ x_1=-1 \text{ and } y_1=4}$$ $$\Huge \text {let }{ x_2=-5 \text{ and } y_2=2}$$ $$\Huge \text{slope point form}$$ $$\Huge y - y_1 = m(x - x_1)$$ $$\Huge{where}$$ $$\Huge m = \frac{(y_2 - y_1)}{(x_2 - x_1)} $$